OpenStudy (anonymous):

How many grams of NaOH are required to make 0.10 moles of NaOH? @Abhisar

OpenStudy (abhisar):

Tell em molecular mass of Na , O and H

OpenStudy (abhisar):

*me

OpenStudy (abhisar):

Na =23, O=16 & H=1

OpenStudy (abhisar):

Molar mass of NaOH= 23+16+1=40

OpenStudy (anonymous):

22.989769280 ± 0.000000020 g/mol Na

OpenStudy (anonymous):

H 1.007940 ± 0.000070 g/mol

OpenStudy (abhisar):

40 grams of NaOH is required for making 1 mole, so for 0.10 moles 40*0.10=0.4 grams of NaOH is required.

OpenStudy (anonymous):

O 15.99940 ± 0.00030 g/mol

OpenStudy (abhisar):

got it ?

OpenStudy (anonymous):

Yea Ok Thanks so much

OpenStudy (abhisar):

Did you know you can give medals to people who help you with questions? Just click the \(\color{blue}{\text{blue "best response" button}}\) to show appreciation for the people who help you.

OpenStudy (abhisar):

\(\color{red}{\huge\bigstar}\huge\text{You're Most Welcome! }\color{red}\bigstar\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)

OpenStudy (abhisar):

\(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\color{blue}{\mathfrak{~~~~Welcome~to~OpenStudy!~\ddot\smile}}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

OpenStudy (anonymous):

:)